# SIRD: A Sender-Informed, Receiver-Driven Datacenter Transport Protocol

### v4

@@ Abstract.
Datacenter congestion management protocols must navigate the throughput-latency-buffering trade-off in the presence of growing constraints due to switching hardware trends, oversubscribed topologies, and varying network configurability and features. In this context, receiver-driven protocols, which schedule packet transmissions instead of reacting to congestion, have shown great promise and work exceptionally well when the bottleneck lies at the ToR-to-receiver link. However, independent receiver schedules may collide if a shared link is the bottleneck instead.
We present SIRD, a receiver-driven congestion control protocol designed around the simple insight that single-owner links should be scheduled while shared links should be managed through traditional congestion control algorithms. The approach achieves the best of both worlds by allowing precise control of the most common bottleneck and robust bandwidth sharing for shared bottlenecks. SIRD is implemented by end hosts and does not depend on Ethernet priorities or extensive network configuration.
We compare SIRD to state-of-the-art receiver-driven protocols (Homa, dcPIM, and ExpressPass) and production-grade reactive protocols (Swift and DCTCP) and show that SIRD is the only one that can consistently maximize link utilization, minimize queuing, and obtain near-optimal latency across a wide set of workloads and traffic patterns. SIRD causes 12×12$\times$12 × less peak buffering than Homa and achieves competitive latency and utilization without requiring Ethernet priorities. Unlike dcPIM, SIRD operates without latency-inducing message exchange rounds and outperforms it in utilization, buffering, and tail latency by 9%, 43%, and 46% respectively. Finally, SIRD achieves 10×10$\times$10 × lower tail latency and $26$% higher utilization than ExpressPass.
copyright: noneconference: ACM Conference; 2024/06;
1. Introduction
@@
Emerging datacenter workloads like ML training (Rajasekaran et al., 2023; Gebara et al., 2021) and disaggregated resource management (Gao et al., 2016) increasingly demand high-throughput and low-latency networking. At the same time, datacenter networking hardware continues to offer higher link speeds, but switch packet buffer capacity is failing to keep up, and SRAM density trends show that this is unlikely to change (Chang et al., 2022; David Schor, 2022). The combination imposes a challenge for congestion control that generally relies on buffering to deliver high throughput. The challenge is compounded by other important concerns such as multi-tenancy, hardware heterogeneity, and cost of operations (Greenberg et al., 2009). In industry, the Ultra Ethernet Consortium is in the process of drafting modern congestion control mechanisms (Consortium, 2023).
@@
The most established approach to managing congestion is by reactively slowing down sender transmission rates after detecting a problem. This is the approach of sender-driven (SD) protocols (Li et al., 2019; Alizadeh et al., 2010; Kumar et al., 2020; Addanki et al., 2022; Zhu et al., 2015; Mittal et al., 2015; Alizadeh et al., 2012; Liu et al., 2021; Arslan et al., 2023; Wang et al., 2023; Zhou et al., 2021) like DCTCP (Alizadeh et al., 2010) and Swift (Kumar et al., 2020) in which senders make decisions based on network feedback. The protocol family has seen decades of evolution in terms of congestion algorithm and signal quality, contributed both by industry and academia. However, the reactive nature of SD protocols often requires several round-trips to address congestion; a limitation for workloads dominated by small flows (Montazeri et al., 2018; Cai et al., 2022; Ghabashneh et al., 2022). Further, the flow-oriented basis of these schemes makes it difficult to reduce latency through message-centric scheduling (Alizadeh et al., 2013; Bai et al., 2015).
@@
To address these limitations, a recent line of work (Hu et al., 2020; Montazeri et al., 2018; Cho et al., 2017; Handley et al., 2017; Cai et al., 2022; Perry et al., 2014; Gao et al., 2015) proactively schedules packet transmissions instead of reacting to congestion buildup. Scheduling is usually handled by receivers that precisely allocate bandwidth by controlling the arrival rate of credit packets to senders. This enables tight control over ToR-to-Host ports which are the most congested (Singh et al., 2015; Zhang et al., 2017; Ghabashneh et al., 2022). The receiver-driven (RD) approach enables high throughput with limited buffering in dcPIM (Cai et al., 2022) and near-optimal latency in Homa (Montazeri et al., 2018).
@@
The fundamental tension in the design of receiver-driven schemes is how to schedule packets over shared links. Each receiver has exclusive control of its downlink but must share the network core and sender uplinks with other receivers. In a distributed protocol, this creates scheduling conflicts if receivers do not coordinate in some manner. Proposed solutions for this problem include explicit pre-matching of senders and receivers (dcPIM (Cai et al., 2022)), in-network credit throttling (Expresspass (Cho et al., 2017)), and the overcommitment of receiver downlinks (Homa (Montazeri et al., 2018)). Despite impressive results, each of these approaches makes a major sacrifice either in message latency, protocol complexity or packet buffer utilization.
We propose SIRD, a receiver-driven design based on the following simple insight: exclusive links (receiver downlinks) should be managed proactively, and shared links (sender uplinks and switch-to-switch links) should be managed reactively. We implement this principle through receiver-driven scheduling that reacts to congestion feedback from shared links using end-host signals to detect congestion on sender uplinks and ECN (Ramakrishnan et al., 2001) to signal congestion in the network. Our approach maintains explicit control over the most common bottleneck (ToR-host links) and simultaneously inherits the robustness, generality, and simplicity of reactive designs. We show that SIRD matches or exceeds the best result of any existing proposal without suffering any of the drawbacks. Beyond quantifiable performance metrics, SIRD’s end-to-end design does not require the use of in-network priority queues, typically reserved to offer isolation among services (Arslan et al., 2023), nor extensive switch configuration.
SIRD receivers coordinate to share core and sender links through a reactive control loop similar to SD schemes. Congestion information flows to receivers within data packets. Receivers run two independent control loops to converge to a bandwidth allocation that does not cause congestion: one to detect congested end-host senders, and one to control congestion in the core network. As network queuing is kept at a minimum, and because senders can initiate data transfers without waiting for matching roundtrips, SIRD delivers messages with low latency.
@@
We compare SIRD to three state-of-the-art proactive protocols (Homa (Montazeri et al., 2018), dcPIM (Cai et al., 2022), and ExpressPass (Cho et al., 2017)) and two production-grade reactive protocols (DCTCP (Alizadeh et al., 2010) and Swift (Kumar et al., 2020)) using large-scale simulations across nine combinations of workloads and traffic patterns. Our evaluation shows that SIRD is the only protocol that can simultaneously maximize link utilization, minimize the use of switch buffers, and obtain near-optimal latency. Specifically: (1) SIRD causes 12×12$\times$12 × less peak top-of-rack (ToR) buffering than Homa, yet achieves competitive latency and utilization in all scenarios without requiring Ethernet priorities. (2) Unlike dcPIM, SIRD does not rely on latency-inducing message exchange rounds and outperforms it in link utilization, peak ToR buffering, and tail latency by 9%, 43%, and 46% respectively while delivering more predictable performance. (3) SIRD achieves 10×10$\times$10 × lower tail latency and $26$% higher utilization than ExpressPass. (4) SIRD outperforms DCTCP and Swift across the board, especially in incast-heavy scenarios.
We further show through a prototype DPDK implementation that SIRD can be implemented at line rate on commodity hardware.
All artifacts will be open-sourced prior to publication.
2. Background
@@
The main goals of congestion control (CC) are enabling efficient sharing of network resources and achieving high network utilization while preventing packet loss. There are also other, at times conflicting, CC goals, e.g., fairness, flow completion time, flow deadlines, and multi-tenancy.
@@
Existing datacenter CC schemes for Ethernet networks can be broadly split into two categories: reactive and proactive. In reactive (also sender-driven (SD)) schemes (Alizadeh et al., 2010, 2012, 2013; Liu et al., 2021; Kumar et al., 2020; Li et al., 2019; Mittal et al., 2015; Zhu et al., 2015), senders control congestion by participating in a distributed coordination process to determine the appropriate transmission rates. Senders obtain congestion information through congestion signals like ECN (Ramakrishnan et al., 2001; Alizadeh et al., 2010), delay (Mittal et al., 2015; Kumar et al., 2020), and in-network telemetry (Li et al., 2019; Wang et al., 2023).
@@
In proactive schemes (Montazeri et al., 2018; Gao et al., 2015; Bai et al., 2015; Handley et al., 2017; Cho et al., 2017; Cai et al., 2022; Perry et al., 2014; Le et al., 2021) bandwidth is allocated explicitly, either globally or per-receiver. Global approaches (Perry et al., 2014; Le et al., 2021) use a centralized arbiter that controls all transmissions and hence face scalability challenges. In Receiver-driven (RD) protocols, each receiver explicitly schedules its downlink by transmitting special-purpose packets such as credits (Cho et al., 2017), grants (Montazeri et al., 2018), or pulls (Handley et al., 2017) to senders. The credit rate can be explicitly controlled by a pacer (Cho et al., 2017; Handley et al., 2017) or be self-clocked (Montazeri et al., 2018; Gao et al., 2015; Cai et al., 2022).
2.1. Exclusive and shared links
@@
Congestion can occur in Host-to-ToR uplinks and the network core, which are shared among receivers, and ToR-to-host downlinks, which are exclusive to a single receiver.
@@
Exclusive links: In the context of RD protocols, ToR-to-Host links (downlinks) can be seen as exclusively controlled by a single entity, the receiver. By controlling the rate of credit transmission, a receiver can explicitly control the rate of data arrival - assuming the bottleneck is the downlink. In fact, downlinks are the most common point of congestion in datacenter networks (Singh et al., 2015; Zhang et al., 2017; Kumar et al., 2020; Ghabashneh et al., 2022). Congestion at the ToR downlink is the result of incast traffic from multiple senders to one receiver. In turn, incast is the result of the fan-out/fan-in patterns of datacenter applications (Dean and Barroso, 2013).
@@
RD schemes excel in managing incast traffic because each receiver explicitly controls the arrival rate of data (Montazeri et al., 2018; Gao et al., 2015; Bai et al., 2015; Handley et al., 2017; Cho et al., 2017; Cai et al., 2022). This level of control also allows RD schemes to precisely dictate which message should be prioritized in downlinks, and can even factor-in a server’s application requirements (Ousterhout et al., 2019). Recent work has delivered significant message latency gains by scheduling messages based on their remaining size (SRPT policy) (Hu et al., 2020; Montazeri et al., 2018; Gao et al., 2015). In contrast, SD schemes (Alizadeh et al., 2010, 2012, 2013; Liu et al., 2021; Kumar et al., 2020; Li et al., 2019; Mittal et al., 2015) treat downlinks as any other link. Senders must first detect downlink congestion and then independently adjust their rates/windows such that the level of congestion falls below an acceptable target.
@@
Shared links: Unlike exclusive downlinks, core and sender-to-ToR link bottlenecks pose a challenge for RD schemes. Shared link bottlenecks can appear when multiple receivers concurrently send credit packets through the same link to pull data in the reverse direction. Before discussing existing approaches to deal with this key challenge, we delve deeper into the specifics of shared-link congestion.
@@
When core congestion occurs, multiple flows with unrelated senders and receivers compete for bandwidth at the network core. Note that the core may have multiple tiers or be configured differently but is fundamentally shared infrastructure. Congestion at the core of a fabric is less common than at downlinks (Montazeri et al., 2018; Singh et al., 2015; Ghabashneh et al., 2022) but can still occur due to core network oversubscription. Oversubscription can be permanent to reduce cost (Singh et al., 2015) or transient due to component failures. Congestion at the core can also occur as a consequence of static ECMP IP routing decisions that cause multiple flows to saturate one core switch and one core-to-Tor downlink, while other core switches have idle capacity (Greenberg et al., 2009; Alizadeh et al., 2014).
@@
Uplink congestion occurs due to fan-out of multiple flows to different receivers or due to bandwidth mismatch between the sender uplink and the receiver downlink. Because the resulting packet buffering is in hosts and not in the fabric, sender congestion is generally a less severe problem. For RD schemes, uplink congestion is known as the unresponsive sender problem and leads to degraded throughput as independent scheduling decisions of receivers may conflict, wasting downlink bandwidth. Whereas the term unresponsive has been used in the context of SRPT scheduling, where messages are transmitted to completion, we will use congested sender as a general description of uplink congestion.
@@
RD protocols have proposed various mechanisms for overcoming the tension between independent receiver scheduling and the fact that some links are shared. One of the earliest designs, pHost (Gao et al., 2015), employs a timeout mechanism at receivers to detect unresponsive senders and direct credit to other senders. NDP (Handley et al., 2017) employs custom very shallow buffer switches and eagerly drops packets, using header trimming to recover quickly. NDP receivers handle uplink sharing by only crediting senders as long as they transmit data packets - whether dropped or not. pHost and NDP do not explicitly deal with core congestion and, further, the analysis by Montazeri et al. (Montazeri et al., 2018) showed that neither of the two achieves high overall link utilization. Homa (Montazeri et al., 2018) introduced controlled overcommitment in which each receiver can send credit to up to $k$ senders at a time. Homa achieves high utilization as it is statistically likely that at least one of the k senders will respond. However, it trades queuing for throughput by meaningfully increasing the amount of expected inbound traffic to each receiver. To let small messages bypass long network queues, Homa relies on Ethernet priorities which are normally used for application-level QoS guarantees (Arslan et al., 2023; Kumar et al., 2020; Hoefler et al., 2023). dcPIM (Cai et al., 2022) employs a semi-synchronous round-based matching algorithm where senders and receivers exchange messages to achieve a bipartite matching. This coordinates the sharing of sender uplinks but congestion at the core is only implicitly addressed by the protocol’s overall low link contention. The downside of this link-sharing approach is message latency. dcPIM delivers small messages quickly by excluding them from the matching process. However, messages larger than the bandwidth-delay-product (BDP) of the network must wait for several RTTs before starting transmission. ExpressPass (Cho et al., 2017) manages all links, exclusive and shared, via a hop-by-hop approach which configures switches to drop excess credit packets, which in turn rate limits data packets in the opposite direction. To reduce credit drops, ExpressPass uses recent credit drop rates as feedback to adjust the future sending rate, and in this way also improves utilization and fairness across multiple bottlenecks. Out of the designs discussed so far, ExpressPass is the only one that explicitly manages core congestion and can operate with highly oversubscribed topologies. ExpressPass’s hop-by-hop design helps it achieve near-zero queuing (Cho et al., 2017), but is more complex to deploy and maintain due to its switch configuration and path symmetry requirements, and suffers under small-flow-dominated workloads (see §5.3).
2.2. The impact of ASIC trends on buffering
Figure 1. Packet buffer size divided by bisection bandwidth for switch ASICs released since 2010 by Broadcom (Xs) and nVidia (circles). See Appendix A for the dataset.

### v5

@@ Abstract
Datacenter congestion control protocols are challenged to navigate the throughput-buffering trade-off while relative packet buffer capacity is trending lower year-over-year. In this context, receiver-driven protocols — which schedule packet transmissions instead of reacting to congestion — excel when the bottleneck lies at the ToR-to-receiver link. However, when multiple receivers must use a shared link (e.g., ToR to Spine), their independent schedules can conflict.
We present SIRD, a receiver-driven congestion control protocol designed around the simple insight that single-owner links should be scheduled, while shared links should be managed with reactive control algorithms. The approach allows receivers to both precisely schedule their downlinks and to coordinate over shared bottlenecks. Critically, SIRD also treats sender uplinks as shared links, enabling the flow of congestion feedback from senders to receivers, which then adapt their scheduling to each sender’s real-time capacity. This results in tight scheduling, enabling high bandwidth utilization with little contention, and thus minimal latency-inducing buffering in the fabric.
We implement SIRD on top of the Caladan stack and show that SIRD’s asymmetric design can deliver 100Gbps in software while keeping network queuing minimal. We further compare SIRD to state-of-the-art receiver-driven protocols (Homa, dcPIM, and ExpressPass) and production-grade reactive protocols (Swift and DCTCP) and show that SIRD is uniquely able to simultaneously maximize link utilization, minimize queuing, and obtain near-optimal latency.
1 Introduction
@@
Datacenter workloads like ML training [62, 32] and disaggregated resource management [30] increasingly demand high-throughput and low-latency networking. While datacenter hardware continues to offer higher link speeds, switch packet buffer capacity is failing to keep up, and SRAM density trends show that this is unlikely to change [17, 24]. The combination imposes a challenge for congestion control, which generally relies on buffering to deliver high throughput. The challenge is compounded by other important concerns such as hardware heterogeneity and cost of operations [35, 23].
@@
The most established approach to managing congestion is by reactively slowing down sender transmission rates after detecting a problem. This is the approach of sender-driven (SD) protocols [47, 8, 44, 4, 74, 54, 9, 50, 12, 69, 73, 67] like DCTCP [8] and Swift [44] in which senders make decisions based on network feedback. The reactive nature of SD protocols often requires several roundtrips to address congestion; a limitation for workloads dominated by small flows [56, 16, 33]. Further, the flow abstraction makes it difficult to reduce latency through message-centric scheduling [10, 13].
@@
To address these limitations, a recent line of work [39, 56, 20, 36, 16, 61, 31] proactively schedules packet transmissions instead of reacting to congestion buildup. Scheduling is typically implemented by having receivers send credit packets to senders, which the latter consume to send back data. This enables tight control over ToR-to-Host ports which are usually the most congested [65, 72, 33]. The receiver-driven (RD) approach enables high throughput with limited buffering in dcPIM [16] and near-optimal latency in Homa [56].
@@
The fundamental tension in the design of receiver-driven schemes is how to schedule packets over shared links. Each receiver has exclusive control of its downlink but must share the bandwidth of the network’s core and that of sender uplinks with other receivers. In a distributed protocol, this creates scheduling conflicts if receivers do not coordinate. Proposed solutions for this problem include explicit pre-matching of senders and receivers (dcPIM [16]), in-network credit throttling (ExpressPass [20]), and the overcommitment of receiver downlinks (Homa [56]). Despite impressive results, each of these approaches makes a sacrifice in either message latency, protocol complexity, or packet buffer utilization.
We propose SIRD, a receiver-driven design based on the following simple insight: exclusive links (receiver downlinks) should be managed proactively, and shared links (sender uplinks and switch-to-switch links) should be managed reactively. In practice, this means that receivers gather congestion feedback from shared links which gives them the means to coordinate, while still explicitly controlling their downlinks. By deducing the bandwidth availability of shared links, and especially that of senders, receivers can be more precise in their credit allocation. This allows them to drive high link utilization without inducing contention and the corresponding queuing in the network. Low queuing, in turn, enables fast message delivery as the network does not introduce head-of-line blocking.
SIRD implements receiver-driven scheduling that reacts to shared-link congestion feedback using 1) end-host signals to detect congestion on sender uplinks, and 2) ECN [63] to detect congestion in the network’s core. Congestion information flows via data packets to receivers and is fed into two independent control loops, with the most congested determining how much credit can be allocated to any given sender. SIRD’s design is end-to-end and does not require the use of in-network priority queues or unconventional switch configuration.
We implement SIRD in 4300 SLOC within Caladan [29], a state-of-the-art dataplane operating system. We show that SIRD can efficiently allocate credit, minimize queuing, and implement round-robin and SRPT policies at 100Gbps.
@@
We use network simulation to systematically compare SIRD to three state-of-the-art proactive protocols (Homa [56], dcPIM [16], and ExpressPass [20]) and two production-grade reactive protocols (DCTCP [8] and Swift [44]). Our evaluation shows that SIRD can simultaneously maximize link utilization, minimize queuing, and obtain near-optimal latency. Specifically: (1) SIRD causes 12×12$\times$12 × less peak top-of-rack (ToR) buffering than Homa yet achieves competitive latency and utilization. (2) Compared to dcPIM, SIRD has no message exchange rounds before sending and outperforms it in link utilization, peak ToR buffering, and tail latency by 9%, 43%, and 46%. (3) SIRD achieves 10×10$\times$10 × lower tail latency and $26$% higher utilization than ExpressPass. (4) SIRD outperforms DCTCP and Swift across the board, especially in incast-heavy scenarios.
2 Background
@@
The main goal of congestion control (CC) is to allocate network resources in a work-conserving manner while avoiding packet loss. There are also other, at times conflicting, CC goals, e.g., fairness, flow completion time or deadlines, and multi-tenancy.
@@
Existing datacenter CC schemes for Ethernet networks can be broadly split into two categories: reactive and proactive. In reactive (also sender-driven (SD)) schemes [8, 9, 10, 50, 44, 47, 54, 74, 67], senders control congestion by participating in a distributed coordination process to determine the appropriate transmission rates. Senders obtain congestion information through congestion signals like ECN [63, 8, 67], delay [54, 44], and in-network telemetry [47, 69].
@@
In proactive schemes [56, 31, 13, 36, 20, 16, 61, 46, 5] bandwidth is allocated explicitly, either globally, or in a distributed manner. Global approaches [61, 46], use a centralized arbiter that controls all transmissions and hence face scalability challenges. Distributed proactive designs operate either end-to-end [56, 31, 13, 16] or with switch involvement in credit management [5, 20]. In end-to-end protocols, which are receiver-driven (RD), each receiver explicitly schedules its downlink by transmitting special-purpose credit tokens to senders. The credit rate can be explicitly controlled by a pacer [20, 36] or be self-clocked [56, 31, 16].
2.1 Exclusive and shared links
@@
Packets flowing to a receiver host can experience congestion either at the ToR-host link, which is exclusive to the receiver, or at host-ToR uplinks and the network core, which are shared among multiple receiving hosts.
@@
Exclusive links: In the context of RD protocols, ToR-to-Host links (downlinks) can be seen as exclusively controlled by a single entity, the receiver. By controlling the rate of credit transmission, a receiver can explicitly control the rate of data arrival - assuming the bottleneck is the downlink. In fact, downlinks are the most common point of congestion in datacenter networks [65, 72, 44, 33]. Congestion at the ToR downlink is the result of incast traffic from multiple senders to one receiver. In turn, incast is the result of the fan-out/fan-in patterns of datacenter applications [25].
@@
RD schemes excel in managing incast traffic because each receiver explicitly controls the arrival rate of data [56, 31, 13, 36, 20, 16]. This level of control also allows RD schemes to precisely dictate which message should be prioritized at downlinks and can even factor-in application requirements [60]. Recent work has delivered significant message latency improvements by scheduling messages based on their remaining size (SRPT policy) [39, 56, 31]. In contrast, SD schemes [8, 9, 10, 50, 44, 47, 54] treat downlinks as any other link. Senders must first detect incast at a receiver’s downlink and then independently adjust their rates/windows such that the level of congestion falls below an acceptable target.
@@
Shared links: Unlike exclusive downlinks, core and sender-ToR link bottlenecks pose a challenge for RD schemes. Shared link bottlenecks can appear when multiple receivers concurrently pull data packets over the same link. Before discussing existing approaches to deal with this key challenge, we delve deeper into the specifics of shared-link congestion.
@@
When congestion occurs in the network core, multiple flows from unrelated senders and receivers compete for bandwidth on core links. Note that although the core may comprise multiple tiers, it is fundamentally shared infrastructure. Congestion at the core of a fabric is less common than at downlinks [56, 65, 33] but can still occur due to core network oversubscription. Oversubscription can be permanent to reduce cost [65] or transient due to component failures. Congestion at the core can also occur because of static ECMP IP routing decisions that cause multiple flows to saturate one core switch and one core-Tor downlink, while other core switches have idle capacity [35, 7].
@@
Uplink congestion occurs due to the fan-out of multiple flows to different receivers or due to bandwidth mismatch between the sender’s uplink and the receiver’s downlink. Because the resulting packet buffering is in hosts and not in the fabric, sender congestion is a less severe problem from the perspective of packet loss. For RD schemes, uplink congestion is known as the unresponsive sender problem [56] and leads to degraded throughput as independent scheduling decisions of receivers may conflict, wasting downlink bandwidth. For example, two receivers may both decide to schedule the same sender $A$ at the same time, even if one of the receivers could be receiving data from sender $B$. Whereas the term unresponsive has been used in the context of SRPT scheduling, where messages are transmitted to completion, we will use congested sender as a general description of uplink congestion.
@@
RD protocols have proposed various mechanisms for overcoming the tension between independent receiver scheduling and the fact that some links are shared. One of the earliest designs, pHost [31], employs a timeout mechanism at receivers to detect unresponsive senders and direct credit to other senders. NDP [36] employs very shallow buffer switches and eagerly drops packets, using packet trimming to recover quickly. NDP receivers handle uplink sharing by only crediting senders if they transmit data packets - whether dropped or not. pHost and NDP do not explicitly deal with core congestion and, further, analysis by Montazeri et al. [56] showed that neither of the two achieves high overall link utilization. Homa [56] introduced controlled overcommitment in which each receiver can send credit to up to $k$ senders at a time. Homa achieves high utilization as it is statistically likely that at least one of the k senders will respond. However, it trades queuing for throughput by meaningfully increasing the amount of expected inbound traffic to each receiver. To let short messages bypass network queues, Homa leverages switch priority queues, which are typically used for application-level QoS guarantees [12, 44, 37]. dcPIM [16] employs a semi-synchronous round-based matching algorithm where senders and receivers exchange messages to achieve a bipartite matching. This coordinates the sharing of sender uplinks but congestion at the core is only implicitly addressed by the protocol’s overall low link contention. The downside of this link-sharing approach is message latency. dcPIM delivers small messages quickly by excluding them from the matching process. However, messages larger than the bandwidth-delay-product (BDP) of the network must wait for several RTTs before starting transmission. ExpressPass [20] manages all links, exclusive and shared, via a hop-by-hop approach which configures switches to drop excess credit packets, which in turn rate limits data packets in the opposite direction. To reduce credit drops, ExpressPass uses recent credit drop rates as feedback to adjust the future sending rate, and in this way also improves utilization and fairness across multiple bottlenecks. Out of the designs discussed so far, ExpressPass is the only one that explicitly manages core congestion and can operate with highly oversubscribed topologies. ExpressPass’s hop-by-hop design helps it achieve near-zero queuing [20] but is more complex to deploy and maintain due to its switch configuration and path symmetry requirements.
2.2 The impact of ASIC trends on buffering